# Quantifying the dynamic structural resilience of international staple food trade networks: An entropy-based approach

### v2

@@ 4.3 Relative contributions of changes in P, PMI, and PMR to changes in resilience
We calculate the relative contributions of changes in the concentration degree of trade flows, economies inter-dependency, and economies inter-independency to changes in network resilience for 1986-2022 (Fig. 6a-d). Overall, changes in the concentration degree of trade flows contribute more to changes in resilience for maize, soybean, and wheat throughout the entire period of 1986-2022, indicating the greater influence of trade flows concentration for food demand on resilience (Liang et al., 2020). However, these effects are diminishing with trade globalization, resulting in negative impact on changes in resilience. Indeed, these findings reflect the underlying shift in food trade discourse from flow monopolies to strategic interactions within relational networks. For example, resilience for rice has been dominated by changes in the concentration degree of trade flows until 2010. Since then, effects of relationships among economies has been highlighted, resulting in decreasing resilience. Although the effect for the concentration degree of trade flows of the rice trade network is higher than 5% after 2010, changes in resilience exhibit phase inversion due to synergistic interaction for interrelationships (Fig. 6b).
@@ 5.1 Relative contributions of changes in economies and trade relationships to changes in resilience
To quantify the relative contributions of changes in economies and trade relationships to changes in resilience, following the leave-one-out approach, we remove some given economy or trade relationship independently and regard changes in the network resilience as the significance of the economy or trade relationship. Figs. 7 and 8 depict the correlation between trade volume of the economy and trade relationship and their contributions to resilience during 1992 to 2022, respectively.
@@
Figure 7: The correlations between trade volume of economy $i$ and changes in resilience of the international food trade network as the economy $i$ is removed during 1992 to 2022. a-f, The maize trade networks. g-l, The rice trade networks. m-r, The soybean trade networks. s-x, The wheat trade networks.
@@
Figure 8: The correlations between trade volume from economy $i$ to $j$ and changes in resilience of the international food trade network as the relationship $e_{ij}$ is removed during 1992 to 2022. a-f, The maize trade networks. g-l, The rice trade networks. m-r, The soybean trade networks. s-x, The wheat trade networks.
@@
Overall, the scale-free characteristics of the international food trade system are reaffirmed from the perspective of changes in resilience caused by economies and trade relationships. That is, a small number of economies and trade relationships contribute most of the changes in resilience, and their influences on resilience demonstrates a power-law distribution. For example, in 1998, only four economies and a single trade relationship contribute more than 2% to resilience in the international wheat trade network, while all remaining economies and relationships each contributes less than 1% (Fig. 7t and Fig. 8t). Over time, the disparity in participants’ contributions to network resilience has narrowed, particularly in the maize and wheat trade networks, indicating that mature networks possess more balanced structures. For example, while four economies and a single trade relationship contribute more than 2.5% to resilience of the international maize trade network in 1992 and 1998 (Fig. 7a,b and Fig. 8a,b), only a single economy contributes more than 2% to that in 2022 (Fig. 7f), as well as all trade relationships contributing less than 1% (Fig. 8f). This finding suggests that more economies and trade relationships are integrating in the system and lessening the significance of the dominators and thus making the system more balanced and resilient, aligning with previous findings (Sartori and Schiavo, 2015).
To further investigate the heterogeneity among leading contributors, we respectively present the top 10 economies and trade relationships in 2022 in Tables LABEL:Table_Economies and LABEL:Table_Relationships, based on their relative contributions to resilience.
@@
Overall, the contributions of leading economies and trade relationships to resilience are highest in the soybean trade network, followed by those in rice, maize, and wheat, which aligns with the resilience characteristics discussed above. It shows that the top 10 economies’ $\Delta R$ values range from 1.64% to 17.06% for soybean, compared to 0.82%-5.79% for rice, 0.62%-2.30% for maize, and 0.34%-1.14% for wheat. Also, the top 10 trade relationships’ $\Delta R$ values range from 1.75% to 23.59% for soybean, 0.70%-2.55% for rice, 0.49%-0.74% for maize, and 0.23%-0.31% for wheat. These findings suggest that the soybean trade network exhibits substantially greater vulnerability to disruptions compared to its maize and wheat counterparts.
@@
Comparison reveals that the contributions of leading economies to network resilience are generally higher than those of primary trade relationships, which can be attributed to the latter’s higher redundancy. For example, while Japan’s maize imports from the United States constitute a critical trade linkage, supply chain disruption in this corridor induces merely a 0.74% impact on the network resilience, owing to substitutable sourcing from Brazil. Reciprocally, Brazilian maize export restrictions would induce a 2.30% impact on the network resilience. The soybean trade network constitutes a notable exception, where individual economy resilience contributions from Brazil (17.06%), China (16.89%), and the United States (13.44%) are eclipsed by the bilateral trade linkages: the United States-China (23.59%) and Brazil-China (20.95%), demonstrating critical path inter-dependency and lack of redundancy.
@@
Although trade volume is significantly positively correlated with the resilience contributions of economies and trade relationships, the two should not be directly equated, particularly when evaluated through the lens of resilience. For a comprehensive comparison, we also present the rankings of economies based on degree centrality and betweenness centrality, two of the most commonly used metrics in the network science literature (Li et al., 2013; Lü et al., 2016; Lambiotte et al., 2019; Liu et al., 2024b), in Table LABEL:Table_Economies. Degree centrality captures the importance of an economy in terms of the number of its trade partners, whereas betweenness centrality reflects its role as a bridge in the network, measured by the number of shortest paths that pass through it. We find that trade volume, degree centrality, and betweenness centrality rankings more accurately capture the strategic importance of economies such as the United States, Germany, France, and Canada. However, all three metrics are fundamentally volume-based and may overlook the dynamic topological structures (Qi and Mei, 2024) and the inter-dependence between nodes (He et al., 2024) within the network. Consequently, these metrics may fail to capture the significance of certain economies such as Paraguay and Kazakhstan with moderate trade volume or low network centrality.
@@
For example, Paraguay ranks sixth (0.75%), eighth (1.46%), and fifth (7.13%) in resilience contributions to the maize, rice, and soybean trade networks, respectively, despite falling outside the top 20 in trade volume and the top 50 in both degree and betweenness centrality for all three commodities. Three potential explanations may account for this phenomenon. First, serving as a strategic redundancy reserve in South America’s multimodal corridors, Paraguay leverages its geopolitical advantage as a landlocked state to activate the Paraguay-Paraná Waterway intermodal system during congestion mitigation protocols at primary Brazilian ports (e.g., Santos). This enables maize rerouting through alternative terminals like Paranaguá, sustaining 15% of displaced shipments while maintaining supply chain continuity. Second, to address the demands of Brazil’s premium culinary market, Paraguay operationalizes a “low-volume and high-premium” trade paradigm by importing Thai Jasmine rice for value-added processing before re-exporting to Brazil. This strategic intervention mitigates production shortfalls in domestic indica rice cultivation. It is further substantiated by the elevated $\Delta R$ values in Paraguay-Brazil and Paraguay-Argentina, as quantified in Table LABEL:Table_Relationships. Third, Paraguay stands as the sole nation maintaining 100% Non-GMO [^8] Non-GMO, Products that are produced without genetic engineering. soybean commercial cultivation globally, positioning itself to fulfill the European Union’s stringent Non-GMO demand (evidenced by a 58% export surge in 2022), thereby securing an irreplaceable supply chain role. Similarly, Kazakhstan ranks fourth (0.64%) in resilience contribution to the wheat trade network, despite falling outside the top 20 in trade volume, degree centrality, and betweenness centrality for this commodity. Its critical role stems from functioning as a crisis replacement corridor, offering alternative routes, such as the China-Europe Railway and the Trans-Caspian Corridor, when the Black Sea passage or conventional land routes are disrupted. Moreover, Kazakhstan supplies 65% of China’s organic wheat imports for animal feed, fulfilling the demand for “Non-GMO and high-gluten” products, thereby securing an indispensable position in the international wheat trade system. Hence, compared to these commonly used methods, we argue that our approach effectively identifies important economies and relationships that possess low volumes but indispensable roles from a resilience perspective. Meanwhile, cases such as Paraguay and Kazakhstan demonstrate that strategic trade positioning may surpass the indiscriminate pursuit of trade volume in enhancing contributions to the international food trade resilience, which may be an indication for other economies.
@@ 5.2 Comparing resilience of core and peripheral food trade networks
Fig. 9 reveals that the observed resilience enhancement directly correlates with efficiency gains, corresponding to the lack of efficiency caused by core trade relationships, aligning with previous findings (Karakoc and Konar, 2021). In other words, overconcentration of flows along core trade pathways can, in fact, reduce both efficiency and resilience of the food trade network. As $p$ increases, resilience of the trade network for maize, rice, soybean, or wheat rises to reach the maximum value ($1/{\mathrm{e}}$) before declining, suggesting that there exists at least one sub-network with optimized resilience, and its metrics are shown in Table 1. Although maize and soybean networks exhibit similar optimal $p$ values, their efficiency effects differ fundamentally. In 2022, for example, the trade-off between efficiency and redundancy in the maize trade network is due to the 68% increase (from 1.23 to 2.07) in efficiency after core relationships are removed. However, the trade-off in the soybean trade network requires the 419% surge (from 0.37 to 1.92) in efficiency, revealing acute oligopolistic vulnerability where the top 2 relationships contribute more than 40% to the network resilience. Also, food trade networks exhibit sharp decline in resilience as extensive trade relationships are removed, while the phenomenon’s $p$ value tends to be closer to 1 in recent years. For example, resilience of the maize trade network decreases sharply (defined as $\Delta R_{p}\leq-\Delta p$ at $p$) after $p=61.41\%$ trade relationships are removed in 1992, while $p=70.43\%$ and $p=91.30\%$ in 2010 and 2022, respectively (Fig. 9c). This finding suggests that the capacity to sustain resilience of peripheral relationships is becoming greater and stronger.
@@
6 Conclusions
Based on the FAO dataset from 1986 to 2022, we construct international trade networks for maize, rice, soybean, and wheat, where economies are represented as nodes and trade relationships among them as directed and weighted links. To address the need of resilience measurement in the networked system, we employ an entropy-based approach grounded in information theory. Subsequently, we explore the correlations between resilience and topological properties. We also employ the index decomposition analysis to quantify the relative contributions of several properties to resilience, including efficiency, redundancy, trade flows concentration, and interrelationship among economies. Furthermore, we apply the leave-one-out approach to identify critical economies and trade relationships from a resilience perspective. Given the pronounced core-periphery structure of networks, we also discuss the roles of core and peripheral components in sustaining network resilience. Our conclusions are presented below.
@@ 6.1 Theoretical contributions
We have several theoretical contributions to the literature on food trade network resilience. First, we introduce a method based on efficiency and redundancy—two fundamental system properties—to measure resilience, drawing from information theory. Our review of existing food trade network studies highlights the need for robust resilience metrics. Unlike prior research that typically relies on changes in unweighted topological properties, we adopt an entropy-based approach originally developed for ecological networks to assess the resilience of the international food trade system. This approach incorporates both the micro perspective of economy and inter-economies behavior and the macro perspective of dynamic network structure, which better focuses on the influence of trade links and flows and effectively measures resilience in food trade systems (addressing RQ1: How can the resilience of the international food trade system can be effectively measured?).
Second, we employ an index decomposition analysis to investigate the contributions of internal components to resilience. Most driving mechanisms in relevant literature focus on topological properties and external shocks, which may overlook the evolution of resilience caused by internal components or structures. To clarify these mechanisms, we decompose resilience into five internal components and investigate their contributions to changes in resilience, including efficiency, redundancy, the concentration degree of trade flows, the inter-dependency among economies, and the inter-independency among economies (addressing RQ2: How do internal components contribute to changes in resilience of the international food trade system?).
Third, by applying a leave-one-out approach, we identify critical economies and trade relationships in the international food trade network from a resilience perspective. While recent literature often employs centrality metrics (e.g., degree centrality and betweenness centrality) to assess the importance of individual economies, probably overlooking critical players with low trade volume or centrality. Moreover, compared to the research on assessing the importance of nodes, that of links are relatively insufficient. Our approach complements the existing literature by capturing critical trade relationships and previously overlooked economies, offering a novel insight compared to traditional metrics (addressing RQ3: How can critical economies and trade relationships in the international food trade system be identified from a resilience perspective?).

### v3

@@ 4.3 Relative contributions of changes in P, PMI, and PMR to changes in resilience
Furthermore, we calculate the relative contributions of changes in the concentration degree of trade flows, economies inter-dependency, and economies inter-independency to changes in network resilience for 1986-2022 (Fig. 6a-d). Overall, changes in the concentration degree of trade flows contribute more to changes in resilience for maize, soybean, and wheat throughout the entire period of 1986-2022, indicating the greater influence of trade flows concentration for food demand on resilience (Liang et al., 2020). However, these effects are diminishing with trade globalization, resulting in negative impact on changes in resilience. Indeed, these findings reflect the underlying shift in food trade discourse from flow monopolies to strategic interactions within relational networks. For example, resilience for rice has been dominated by changes in the concentration degree of trade flows until 2010. Since then, effects of relationships among economies has been highlighted, resulting in decreasing resilience. Although the effect for the concentration degree of trade flows of the rice trade network is higher than 5% after 2010, changes in resilience exhibit phase inversion due to synergistic interaction for interrelationships (Fig. 6b).
@@ 5.1 Relative contributions of changes in economies and trade relationships to changes in resilience
To quantify the relative contributions of changes in economies and trade relationships to changes in resilience, following the leave-one-out approach, we remove some given economy or trade relationship independently and regard changes in the network resilience ($\Delta R$) as its relative contribution. Figs. 7 and 8 respectively depict the log-log correlations between $\Delta R$ and trade volume of both economy and relationship during 1992 to 2022.
@@
Figure 7: The correlations between trade volume of economy $i$ and changes in resilience of the international food trade network as the economy $i$ is removed during 1992 to 2022. In this figure, the red dashed line denotes the fit to scatters, where $k$ is slope and $R^{2}$ is goodness of fit. a-f, The maize trade networks. g-l, The rice trade networks. m-r, The soybean trade networks. s-x, The wheat trade networks.
Overall, trade volume is significantly positively correlated with contributions to resilience, indicating that economies or trade relationships with larger volumes are generally more critical for sustaining network resilience. For economies, however, we find that the slopes in recent years are significantly smaller than those in earlier years. For rice, for example, the slopes exceeded 5.2 during 1992-2010 but decreased to 3.3 after 2016. By contrast, the slopes for trade relationships have remained relatively stable and even increased slightly. These results suggest that the influence of economies’ trade volumes on their resilience contributions is declining, while the role of interactions is stable and becoming more important compared to flow monopolies, which are also consistent with our decomposition analysis of the relative contributions of changes in P, PMI, and PMR (Fig. 6).
@@
Figure 8: The correlations between trade volume from economy $i$ to $j$ and changes in resilience of the international food trade network as the relationship $\varepsilon_{ij}$ is removed during 1992 to 2022. In this figure, the red dashed line denotes the fit to scatters, where $k$ is slope and $R^{2}$ is goodness of fit. a-f, The maize trade networks. g-l, The rice trade networks. m-r, The soybean trade networks. s-x, The wheat trade networks.
@@
Meanwhile, the scale-free characteristics of the international food trade system are reaffirmed from the perspective of changes in resilience caused by economies and trade relationships. That is, a small number of economies and trade relationships contribute most of the changes in resilience, and their influences on resilience demonstrates a power-law distribution. For example, in 1998, only four economies and a single trade relationship contribute more than 2% to resilience in the international wheat trade network, while all remaining economies and relationships each contributes less than 1% [^8] The results are available on request due to space limitations.. Over time, the disparity in participants’ contributions to network resilience has narrowed, particularly in the maize and wheat trade networks, indicating that mature networks possess more balanced structures. For example, while four economies and a single trade relationship contribute more than 2.5% to resilience of the international maize trade network in 1992 and 1998, only a single economy contributes more than 2% to that in 2022, as well as all trade relationships contributing less than 1%. This finding suggests that more economies and trade relationships are integrating in the system and lessening the significance of the dominators and thus making the system more balanced and resilient, aligning with previous findings (Sartori and Schiavo, 2015).
@@
To further investigate the heterogeneity among leading contributors, we respectively present the 2022 top 10 economies and trade relationships in Tables LABEL:Table_Economies and LABEL:Table_Relationships, based on their relative contributions to resilience. Overall, the contributions of leading economies and trade relationships to resilience are highest in the soybean trade network, followed by those in rice, maize, and wheat, which aligns with the resilience characteristics discussed above. It shows that the top 10 economies’ $\Delta R$ values range from 1.64% to 17.06% for soybean, compared to 0.82%-5.79% for rice, 0.62%-2.30% for maize, and 0.34%-1.14% for wheat. Also, the top 10 trade relationships’ $\Delta R$ values range from 1.75% to 23.59% for soybean, 0.70%-2.55% for rice, 0.49%-0.74% for maize, and 0.23%-0.31% for wheat. These findings suggest that the soybean trade network exhibits substantially greater vulnerability to disruptions compared to its maize and wheat counterparts.
@@
Although the results exhibits the positive correlations between trade volume and resilience contributions, according to Tables LABEL:Table_Economies and LABEL:Table_Relationships, the two should not be directly equated. In the case of economies, for a comprehensive comparison, we also present their rankings based on degree centrality, betweenness centrality, PageRank centrality, and eigenvector centrality, four of the most commonly used centrality measures in the network science literature (Li et al., 2013; Lü et al., 2016; Lambiotte et al., 2019; Liu et al., 2024b; Gutiérrez-Moya et al., 2021; Zhang and Zhou, 2023) in Table LABEL:Table_Economies. Among them, degree centrality captures the importance of an economy in terms of the number of its trade partners, whereas betweenness centrality reflects its role as a bridge in the network, measured by the number of shortest paths that pass through it. PageRank centrality and eigenvector centrality emphasize the importance of an economy depends on both the number and the importance of its neighboring economies.
We find that the rankings based on trade volume and four centrality measures capture more accurately the strategic importance of economies such as Spain in the maize trade network, the United States in the rice trade network, Germany in the soybean trade network, and Italy in the wheat trade network. These economies are developed and interact with many partners, thereby occupying both core and bridging positions in the networks. However, such measures may fail to capture the role of some developing producers and exporters, including India, Thailand, and the Philippines in the rice trade network, Brazil in the soybean trade network, and Ukraine in the wheat trade network. These economies are evidently important due to their export capacity and grain yields, yet their trade partners are typically consumers rather than re-exporters. As a result, PageRank and eigenvector centrality often classify them as peripheral nodes, thereby overlooking their strategic importance. Consequently, we propose that $\Delta R$ can complement some conventional centrality measures, as illustrated by the following two cases.
@@
Paraguay ranks sixth (0.75%), eighth (1.46%), and fifth (7.13%) in contributions to resilience for maize, rice, and soybean, respectively, despite falling outside the top 20 in trade volume and the top 50 in four centrality measures for all three commodities. Three potential explanations may account for this phenomenon. First, serving as a strategic redundancy reserve in South America’s multimodal corridors, Paraguay leverages its geopolitical advantage as a landlocked state to activate the Paraguay-Paraná Waterway intermodal system during congestion mitigation protocols at primary Brazilian ports (e.g., Santos). This enables maize rerouting through alternative terminals like Paranaguá, sustaining 15% of displaced shipments while maintaining supply chain continuity (see results of $\Delta R=0.62\%$ for Paraguay-Brazil maize trade relationship in Table LABEL:Table_Relationships). Second, to address the demands of Brazil’s premium culinary market, Paraguay operationalizes a “low-volume and high-premium” trade paradigm by importing Thai Jasmine rice for value-added processing before re-exporting to Brazil. This strategic intervention mitigates production shortfalls in domestic indica rice cultivation. It is further substantiated by the elevated $\Delta R$ values in Paraguay-Brazil and Paraguay-Argentina, as quantified in Table LABEL:Table_Relationships. Third, Paraguay stands as the sole nation maintaining 100% Non-GMO [^9] Non-GMO, Products that are produced without genetic engineering. soybean commercial cultivation globally, positioning itself to fulfill the European Union’s stringent Non-GMO demand (evidenced by a 58% export surge in 2022), thereby securing an irreplaceable supply chain role.
Similarly, Kazakhstan ranks fourth (0.64%) in contributions to resilience for wheat, despite falling outside the top 20 in trade volume and all centrality measures. Its critical role stems from functioning as a crisis replacement corridor, offering alternative routes, such as the China-Europe Railway and the Trans-Caspian Corridor, when the Black Sea passage or conventional land routes are disrupted (e.g., the Russia-Ukraine conflict). Moreover, Kazakhstan supplies 65% of China’s organic wheat imports for animal feed, fulfilling the demand for “Non-GMO and high-gluten” products, thereby securing an indispensable position in the international wheat trade system.
Cases such as Paraguay and Kazakhstan demonstrate that strategic trade positioning may surpass the indiscriminate pursuit of trade volume in enhancing contributions to the international food trade resilience, which may be an indication for other economies.
@@ 5.2 Comparing resilience of core and peripheral food trade networks
Fig. 9 reveals that the observed resilience enhancement directly correlates with efficiency gains, corresponding to the lack of efficiency caused by core trade relationships, aligning with previous findings (Karakoc and Konar, 2021). As $p$ increases, resilience of the trade network for maize, rice, soybean, or wheat rises to reach the maximum value ($1/{\mathrm{e}}$) before declining, suggesting that there exists at least one sub-network with optimized resilience, and its metrics are shown in Table 1. Although maize and soybean networks exhibit similar optimal $p$ values, their efficiency effects differ fundamentally. In 2022, for example, the trade-off between efficiency and redundancy in the maize trade network is due to the 68% increase (from 1.23 to 2.07) in efficiency after core relationships are removed. However, the trade-off in the soybean trade network requires the 419% surge (from 0.37 to 1.92) in efficiency, revealing acute oligopolistic vulnerability where the top 2 relationships contribute more than 40% to the network resilience. Also, food trade networks exhibit sharp decline in resilience as extensive trade relationships are removed, while the phenomenon’s $p$ value tends to be closer to 1 in recent years. For example, resilience of the maize trade network decreases sharply (defined as $\Delta R_{p}\leq-\Delta p$ at $p$) after $p=61.41\%$ trade relationships are removed in 1992, while $p=70.43\%$ and $p=91.30\%$ in 2010 and 2022, respectively (Fig. 9c). This finding suggests that the capacity to sustain resilience of peripheral relationships is becoming greater and stronger.
@@
6 Discussions
@@ 6.1 Theoretical contributions
We have several theoretical contributions to the literature on food trade network resilience. First, our research advances the conceptual foundation of food trade system resilience by demonstrating that it can be understood as the outcome of a trade-off between efficiency and redundancy. Inspired by the assumption that long-term natural selection drives ecosystems to optimize resilience (Ulanowicz, 2009; Liang et al., 2020; Kharrazi et al., 2013), we employ an entropy-based approach to address the first research question. Aligning with Kharrazi et al. (2017), our analysis also suggests that systems that are either over-redundant or over-efficient are not sufficiently resilient because the former wastes excessive resources while the latter is too vulnerable during disruptions (see special cases in Appendix A3 and Appendix A4). Our research not only supports the current view that resilience emerges from balancing multiple principles (Essuman et al., 2020; Brede and de Vries, 2009; Kamalahmadi et al., 2022; Bullock et al., 2017), but also broadens the scope of resilience measurement in international food trade networks and provides an information-theoretic framework for assessing resilience across various trade networks.
Second, our research enriches resilience theory by revealing the internal mechanisms that drive resilience dynamics. By employing an index decomposition analysis (Ang and Zhang, 2000), we explore how resilience evolves through the concentration of trade flows and both inter-dependence and inter-independence among economies and thereby address the second research question, in line with studies on elemental cycling networks (Liang et al., 2020; Luo et al., 2024). Consistent with the existing literature (Li et al., 2024), we also examine the correlations between topological indicators and resilience. Thus, our research extends previous studies that focused mainly on external shocks or network indicators by exhibiting how resilience is shaped by internal components, thereby advancing the understanding of the structural underpinnings of resilience.
Third, our research reexamines the roles of economies and trade relationships in sustaining network resilience. Drawing on the systemic risk literature (Hué et al., 2019; Zedda and Cannas, 2020), we define an economy or relationship as critical if its removal substantially alters the network’s trade-off. While previous research has often emphasized the influence of core economies (Gutiérrez-Moya et al., 2021), our findings identify that peripheral economies, through specific trade linkages, can also play important roles in maintaining resilience, which addresses the third research question. This broadens the theoretical perspective toward a more distributed and inclusive understanding of trade networks, particularly highlighting the significance of peripheral nodes. Moreover, different centrality measures, grounded in distinct assumptions (Borgatti, 2005; Sciarra et al., 2018), may yield divergent results. Our resilience-based perspective complements existing centrality measures and, when applied in conjunction with traditional measures, facilitates a more precise identification of nodes or links that are critical to the network structure.